Directions: Establish the table of information and formulate the LP Model of the following:

1.) Rina needs at least 48 units of protein, 60 units of carbohydrates, and 50 units of fat each month. From each kilogram of food A, she receives 2 units of protein, 4 units of carbohydrates, and 5 units of fats. Food b contains 3 units of protein, 3 units of carbohydrates, and 2 units of fats. If food A costs Php110 per kilogram and food B costs Php 90 per kilogram. How many kilograms of each food should Sofia buy each month to keep costs at a minimum?



1
Expert's answer
2021-11-21T13:31:37-0500

for x kg of A and y kg of B, you want to minimize

p = 100x1 + 90x2

subject to the conditions that

"2x_1+3x_2 \ge 46"

"4x_1+3x_2 \ge 60"

"5x_1+2x_2 \ge 50"


the optimal solution to the given LP problem is : x1=7,x2=10.67 and min z=1660.